11-55. A purchased machine cost $320,000 with delivery and installation charges amounting to $30,000. The declared salvage value was $50,000. Early in year 3, the company changed its product mix and found that it no longer needed the machine. One of its competitors agreed to purchase the machine for $180,000. Determine the loss, gain, or recapture for MACRS depreciation on the sale. The MACRS property class for the machine is 7 years
Answer #1

cost basis = 320000 + 30000 = 350000

Depreciation in yr1 = 0.1429 * 350000 = 50015

Depreciation in yr2 = 0.2449 * 350000 = 85715

Depreciation in yr3 = 0.1749 * 350000 = 61215

Book value in yr 3 = 350000 - 50015 - 85715 - 61215/2 = 183662.50

Gain/Loss after selling machine = 180000 - 183662.50 = -3662.50

Therefore loss of 3662.50